# Mutual-visibility coloring of graphs

### v1

@@ 3 Mutual-visibility coloring in glued $t$-ary trees
Throughout this section, the vertices of two copies of the binary tree in $GT(r)$ are denoted by $v_{i,j}$ and $v_{i,j}^{\prime}$, respectively, for $i=1,2,\ldots,r-1$ and i1i-1 is the depth of $v_{i,j}$ and $v^{\prime}_{i,j}$ in $T_{r}^{(1)}$ and $T_{r}^{(2)}$, respectively; and $j$ is the position of the vertex counted from left to right in the drawing. The quasi-leaves, which are leaf vertices of both rooted binary trees, are denoted by $v_{1},v_{2},......v_{2^{r}}$, labeled from left to right in the drawing.
@@
Given two vertices $v_{i,j}\in T^{(1)}_{r}$, $v^{\prime}_{k,l}\in T^{(2)}_{r}$ where $i<k$ and any leaf vertex $v_{a}$ of $T_{v_{k,l}}$, then there exists a unique geodesic path between $v_{i,j}$ and $v^{\prime}_{k,l}$, which passes through $v_{a}$ and each of these geodesic paths passes through $v_{k,l}$ also. Note that here $j$ and $l$ may be equal or distinct. Also, a similar statement holds for $k<i$.
@@
If both quasi-leaves in $C_{i,j}$ are present in $S$, then $|S\cap Q_{i,j}|\leq 1$ and $|S\cap Q_{i,j}^{\prime}|\leq 1$. Also, $S$ cannot simultaneously intersect $Q_{i,j}$ and $Q_{i,j}^{\prime}$ and hence $|S\cap V(C_{i,j})|\leq 3$. If one of the quasi-leaves in $C_{i,j}$ is present in $S$, then again $|S\cap Q_{i,j}|\leq 1$ and $|S\cap Q_{i,j}^{\prime}|\leq 1$ and hence $|{S\cap V(C_{i,j})}|\leq 3$.
Now consider the case where both the quasi-leaves are not present in $S$. Since the quasi-leaves are not present in $S$, then since $|S\cap Q_{i,j}|\leq 2$ and $|S\cap Q_{i,j}^{\prime}|\leq 2$, we have $|{S\cap V(C_{i,j})}|\leq 4$. But, if there are two vertices each from both $Q_{i,j}$ and $Q^{\prime}_{i,j}$, say $v_{i,j},v_{k,l},v^{\prime}_{a,b}$ and $v^{\prime}_{c,d}$, then from the above observations, at least one pair of these vertices will not be mutually-visible, which is a contradiction. Hence, the lemma. ∎
@@
In this case, assign color $1$ to all the quasi-leaves, color $2k$ to all the vertices of depth rkr-k in $V_{r}^{(1)}$ and to the $k^{th}$ vertex in the sequence $\{v^{\prime}_{1,1},v^{\prime}_{2,1},v^{\prime}_{2,2},v^{\prime}_{3,1},$ $v^{\prime}_{3,2},v^{\prime}_{3,3},v^{\prime}_{3,4},\ldots\}$, for $k=1,2,\ldots,r-i$. Similarly, assign color $2k+1$ to all vertices of depth rkr-k in $V_{r}^{(2)}$ and to the $k^{th}$ vertex in the sequence $\{v_{1,1},v_{2,2},v_{2,1},v_{3,4},v_{3,3},v_{3,2},v_{3,1},\ldots\}$, for $k=1,2,\ldots,r-i$. Here all vertices of depth $0,1,\ldots,i-2$ and $r-1,r-2,\ldots,r-(r-i)=i$ are completely colored and vertices of depth i1i-1 are partially colored. Among the vertices of depth i1i-1 in $V_{r}^{(2)}$, $v^{\prime}_{i-1,1},v^{\prime}_{i-1,2},\ldots,v^{\prime}_{i-1,j}$ will be already colored, where $j=r-i-(2^{i-1}-1)\geq 2^{i-2}-1$, (since, $r\geq 2^{i-1}+2^{i-2}+i-2$). Similarly, in $V_{r}^{(1)}$, $v_{i-1,a},v_{i-1,a-1},\ldots,v_{i-1,a-j-1}$ will be already colored, where $a=2^{i-1}$ and $j\geq 2^{i-2}-1$. When $r\geq 2^{i-1}+2^{i-2}+i-1$, we have $j\geq 2^{i-2}$ so that there are fewer than $2^{i-2}$ uncolored vertices of depth i1i-1. In this case, color all the remaining uncolored vertices with color 2(ri)+22(r-i)+2. When $r=2^{i-1}+2^{i-2}+i-2$, color $v_{i-1,j+1}$ with color 1 and all the remaining vertices with color 2(ri)+22(r-i)+2 to obtain a mutual-visibility coloring of $GT(r)$ with 2(ri)+22(r-i)+2 colors.
@@
Sub Case 2: $r=2^{i-1}+i-1$.
Again we proceed with the same coloring pattern till k=rik=r-i. There are $1+2+2^{2}+\ldots+2^{i-1}=2^{i}-1=r-i-1$ vertices in the subtree of depth i1i-1 rooted at $v_{11}^{\prime}$ and all of them are colored using colors $2,4,\ldots,2(r-i)$. Now, assign the color 2(ri)+22(r-i)+2 to the vertices of depth rir-i in $V_{r}^{(1)}$ and the color 2(ri)+32(r-i)+3 to the vertices of depth rir-i in $V_{r}^{(2)}$. This is a mutual-visibility coloring of $GT(r)$ with 2(ri)+32(r-i)+3 colors.
Sub case 3: $2^{i-1}+i\leq r\leq 2^{i-1}+2^{i-2}+i-3$.
The coloring procedure is the same as Case 1, where the only difference is that more than half of the vertices of depth i1i-1 will be uncolored. We can use one color each for these uncolored vertices in $V_{r}^{(1)}$ and $V_{r}^{(2)}$ so that the resulting coloring is a mutual-visibility coloring of $GT(r)$ by Lemma 3.5.
To prove the converse, first we prove that 2(ri)+12(r-i)+1 colors are not enough for a mutual-visibility coloring of $GT(r)$, where $i$ is the unique integer such that $2^{i-1}+i-2\leq r\leq 2^{i}+i-3$. Assume the contrary that there exists a mutual-visibility coloring with 2(ri)+12(r-i)+1 colors.
@@
Now, consider the cycles $C_{t2^{r-2},t2^{r-2}+1}$, for $t=1,3$. These two cycles have 4r124r-12 vertices each which are not part of both the previous cycles considered. Similar to the case above we can find a set of colors $\chi_{2}$ which are used three times either in $C_{2^{r-2},2^{r-2}+1}-(C_{1,2^{r}}\cup C_{2^{r-1},2^{r-1}+1})$ and $C_{2^{r-1}+2^{r-2},2^{r-1}+2^{r-2}+1}-(C_{1,2^{r}}\cup C_{2^{r-1},2^{r-1}+1})$. Also, $\chi_{2}\cap\chi_{1}=\chi_{2}\cap\chi_{0}=\phi$ and $|\chi_{2}|\geq 2(4i-12)$. Now, we consider the four cycles $C_{t2^{r-3},t2^{r-3}+1}$, for $t=1,3,5,7$, and proceed as above to get $\chi_{3}$ of cardinality at least 4(4i16)4(4i-16) which do not intersect pairwise with any of $\chi_{0},\chi_{1}$ and $\chi_{2}$. Continue like this to get a collection of distinct colors $\chi_{1},\chi_{2},\ldots,\chi_{i-1}$, where each $\chi_{j}$ is of cardinality at least $2^{i-1}(4i-4j-4)$, $j\in[i-1]$ and $\chi_{0}$ of cardinality at least 4i24i-2. Note that the disjoint sets $\chi_{j}$, $j\in[i-1]$ have cardinality at least $2^{i-1}(4i-4j-4)$ which is an arithmetic-geometric progression with common difference -4 and common ratio 2. Therefore, $\cup_{j=1}^{i-1}\chi_{j}\geq 2^{i+1}-4i$ and together with the 4i24i-2 more distinct colors in $\chi_{0}$, at least $2^{i+1}-2$ colors are used three times to color $GT(r)$. But, we have only 2(ri)+12(r-i)+1 colors and $r\leq 2^{i}+i-2$ so that, $2(r-i)+1\leq 2^{i+1}-3$, a contradiction.
@@
[…]
@@
Proof.
In the coloring in the proof of Theorem 3.6, the vertices in each of the color class is also in general position except for $r=2^{i-1}+2^{i-2}+i-2$, where the color of the quasi-leaves is used to color a non-quasi-leaf vertex. Hence the corollary. ∎
@@
Let $S$ be a mutual-visibility set of $GT(r,t)$. $C_{n,k}$ denote the circle which passes through the quasi-leaves $v_{n}$ and $v_{k}$. Then $|{S\cap V(C_{i,j})}|\leq 3$.
@@
We proceed with the same coloring pattern used in Theorem 4.6 so that we reach a stage where the vertices of depth i1i-1 in $V_{r}^{(2)}$, $v^{\prime}_{i-1,1},v^{\prime}_{i-1,2},\ldots,v^{\prime}_{i-1,j}$ will be already colored, where $j=r-i-(\frac{t^{i-1}-1}{t-1})\geq\frac{t^{i-1}}{2}-1$, (since, $r\geq\frac{t^{i-1}-1}{t-1}+\frac{t^{i-1}}{2}+i-1$). Similarly, in $V_{r}^{(2)}$, $v^{\prime}_{i-1,a},v^{\prime}_{i-1,a-1},\ldots,v^{\prime}_{i-1,a-j-1}$ will be already colored, where $a=t^{i-1}$ and $j\geq\frac{t^{i-1}}{2}-1$. When $r\geq\frac{t^{i-1}-1}{t-1}+\frac{t^{i-1}}{2}+i-1$, we have $j\geq\frac{t^{i-1}}{2}$ so that there are fewer than $\frac{t^{i-1}}{2}$ uncolored vertices of depth i1i-1. In this case, color all the remaining uncolored vertices with color 2(ri)+22(r-i)+2. When $r=\frac{t^{i-1}-1}{t-1}+\frac{t^{i-1}}{2}+i-1$, color $v_{i-1,j+1}$ with color 1 and all the remaining vertices with color 2(ri)+22(r-i)+2 to obtain a mutual-visibility coloring of $GT(r,t)$ with 2(ri)+22(r-i)+2 colors.
@@
Sub case 2: $r=\frac{t^{i-1}-1}{t-1}+i$.
Again we proceed same coloring pattern upto k=rik=r-i . There are $1+t+t^{2}+\ldots+t^{i-2}=\frac{t^{i-1}-1}{t-1}=r-i-1$ vertices in the subtree of depth i1i-1 rooted at $v_{11}^{\prime}$ and all of them are colored using colors $2,4,\ldots,2(r-i)$. Now, assign the color 2(ri)+22(r-i)+2 to the vertices of depth rir-i in $V_{r}^{(1)}$ and the color 2(ri)+32(r-i)+3 to the vertices of depth rir-i in $V_{r}^{(2)}$. This is a mutual-visibility coloring of $GT(r,t)$ with 2(ri)+32(r-i)+3 colors.
Sub case 3: $\frac{t^{i-1}-1}{t-1}+i+1\leq r\leq\frac{t^{i-1}-1}{t-1}+\frac{t^{i-1}}{2}+i-2$.
The coloring procedure is the same as Case 1, where the only difference is that more than half of the vertices of depth i1i-1 will be uncolored. We can use one color each for these uncolored vertices in $V_{r}^{(1)}$ and $V_{r}^{(2)}$ so as to obtain a mutual-visibility coloring of $GT(r,t)$.
@@
Consider the cycle $C_{1,t^{r}}$. By Lemma 3.8, a color can be used at most 3 times to color the vertices of this cycle. Since we have only 2(ri)+12(r-i)+1 colors, and there are $4r$ vertices in $C_{1,t^{r}}$, at least 4r2[2(ri)+1]=4i24r-2[2(r-i)+1]=4i-2 colors must be used at least three times. Let $\chi_{0}$ be the collection of colors which are used three times to color the vertices of $C_{1,t^{r}}$ so that $|\chi_{0}|\geq 4i-2$. Now, consider (t1)(t-1) cycles $C_{mt^{r-1},mt^{r-1}+1}$ where $m\in[t-1]$, note that all the cycles have length $4r$. Let $S_{m}=C_{mt^{r-1},mt^{r-1}+1}-(C_{1,t^{r}}\cup C_{(m-1)t^{r-1},(m-1)t^{r-1}+1})$, $m\in[t-1]$ . Note that there are 4r44r-4 vertices in $S_{m}$, if $m\in[t-2]$ and 4r64r-6, if m=t1m=t-1. Similar to the discussion above, at least 4i64i-6 colors must be used three times to color those vertices in $S_{m}$, $m\in[t-2]$ and at least 4i84i-8 colors must be used three times to color vertices in $S_{t-1}$. Let $\chi_{m}$ denotes the collection of colors which are used three times to color the vertices of $S_{m}$ , $m\in[t-1]$ so that $|\chi_{m}|\geq 4i-6$, where $m\in[t-2]$ and $|\chi_{m}|\geq 4i-8$, when m=t1m=t-1.
@@
Generally in $n^{th}$ step we consider $(t-1)t^{n-1}$ cycles $C_{mt^{r-n},mt^{r-n}+1}$ where $m\in[t^{n}]-t\mathbb{N}$. Note that in the $n^{th}$ step all the cycles have length 4(rn+1)4(r-n+1). Set $S^{0}_{m}=C_{1,2^{r}}$ and define $S_{m}^{n}=C_{mt^{r-n},mt^{r-n}+1}-\{C_{1,t^{r}}\cup S_{m}^{n-1}\cup C_{(m-1)t^{r-n},(m-1)t^{r-n}+1}\}$, for every $n\in[i-1]$ . Note that there are 4r4n4r-4n vertices in $S^{n}_{m}$, if $m\ncong 1\pmod{t}$ and 4r4n24r-4n-2, if $m\cong 1\pmod{t}$. Similar to the discussion above, since we have only 2(ri)+12(r-i)+1 colors, at least 4r4n2[2(ri)+1]=4i4n24r-4n-2[2(r-i)+1]=4i-4n-2 colors must be used at least three times to color those vertices in $S_{m}^{n}$, where $m\ncong 1\pmod{t}$, and at least 4i4n44i-4n-4 colors must be used three times to color those vertices in $S_{m}^{n}$, where $m\cong 1\pmod{t}$. Let $\chi^{n}_{m}$ denotes the collection of colors which are used three times to color the vertices of $S_{m}^{n}$, $m\in[t-1]-t\mathbb{N}$ so that $|\chi^{n}_{m}|\geq 4i-4n-2$, if $m\ncong 1\pmod{t}$ and $|\chi^{n}_{m}|\geq 4i-4n-4$, if $m\cong 1\pmod{t}$.

### v2

@@ 3 Mutual-visibility coloring in glued $t$-ary trees
Throughout this section, the vertices of two copies of the binary tree in $GT(r)$ are denoted by $v_{i,j}$ and $v_{i,j}^{\prime}$, respectively, for $i=1,2,\ldots,r-1$ and i1i-1 is the depth of $v_{i,j}$ and $v^{\prime}_{i,j}$ in $T_{r}^{(1)}$ and $T_{r}^{(2)}$, respectively; and $j$ is the position of the vertex counted from left to right in the drawing. The quasi-leaves, which are leaf vertices of both rooted binary trees, are denoted by $v_{1},v_{2},......v_{2^{r}}$, labeled from left to right in the drawing. See Figure 3 for the denotation of the vertices in graph $GT(3)$ and graph $GT(4)$.
@@
Given two vertices $v_{i,j}\in T^{(1)}_{r}$, $v^{\prime}_{k,l}\in T^{(2)}_{r}$ where $i<k$ and any leaf vertex $v_{a}$ of $T_{v_{k,l}}$, then there exists a unique geodesic path between $v_{i,j}$ and $v^{\prime}_{k,l}$, which passes through $v_{a}$ and each of these geodesic paths passes through $v_{k,l}$ also. Also, a similar statement holds for $k<i$.
@@
If both quasi-leaves in $C_{i,j}$ are present in $S$, then $|S\cap V(Q_{i,j})|\leq 1$ and $|S\cap V(Q_{i,j}^{\prime})|\leq 1$. Also, $S$ cannot simultaneously intersect $V(Q_{i,j})$ and $V(Q_{i,j}^{\prime})$ and hence $|S\cap V(C_{i,j})|\leq 3$. If one of the quasi-leaves in $C_{i,j}$ is present in $S$, then again $|S\cap V(Q_{i,j})|\leq 1$ and $|S\cap V(Q_{i,j}^{\prime})|\leq 1$, and hence $|{S\cap V(C_{i,j})}|\leq 3$.
Now consider the case where both the quasi-leaves are not present in $S$. Since the quasi-leaves are not present in $S$, then since $|S\cap V(Q_{i,j})|\leq 2$ and $|S\cap V(Q_{i,j}^{\prime})|\leq 2$, we have $|{S\cap V(C_{i,j})}|\leq 4$. But, if there are two vertices each from both $Q_{i,j}$ and $Q^{\prime}_{i,j}$, say $v_{i,j},v_{k,l},v^{\prime}_{a,b}$ and $v^{\prime}_{c,d}$, then from the above observations, at least one pair of these vertices will not be mutually-visible, which is a contradiction. Hence, the lemma. ∎
@@
In this case, assign color $1$ to all the quasi-leaves, color $2k$ to all the vertices of depth rkr-k in $V_{r}^{(1)}$ and to the $k^{th}$ vertex in the sequence $\{v^{\prime}_{1,1},v^{\prime}_{2,1},v^{\prime}_{2,2},v^{\prime}_{3,1},$ $v^{\prime}_{3,2},v^{\prime}_{3,3},v^{\prime}_{3,4},\ldots\}$, for $k=1,2,\ldots,r-i$. Similarly, assign color $2k+1$ to all vertices of depth rkr-k in $V_{r}^{(2)}$ and to the $k^{th}$ vertex in the sequence $\{v_{1,1},v_{2,2},v_{2,1},v_{3,4},v_{3,3},v_{3,2},v_{3,1},\ldots\}$, for $k=1,2,\ldots,r-i$. Here all vertices of depth $0,1,\ldots,i-2$ and $r-1,r-2,\ldots,r-(r-i)=i$ are completely colored and vertices of depth i1i-1 are partially colored. Among the vertices of depth i1i-1 in $V_{r}^{(2)}$, $v^{\prime}_{i,1},v^{\prime}_{i,2},\ldots,v^{\prime}_{i,j}$ will be already colored, where $j=r-i-(2^{i-1}-1)\geq 2^{i-2}-1$, (since, $r\geq 2^{i-1}+2^{i-2}+i-2$). Similarly, in $V_{r}^{(1)}$, $v_{i,a},v_{i,a-1},\ldots,v_{i,a-j-1}$ will be already colored, where $a=2^{i-1}$ and $j\geq 2^{i-2}-1$. When $r\geq 2^{i-1}+2^{i-2}+i-1$, we have $j\geq 2^{i-2}$ so that there are fewer than $2^{i-2}$ uncolored vertices of depth i1i-1. In this case, color all the remaining uncolored vertices with color 2(ri)+22(r-i)+2. When $r=2^{i-1}+2^{i-2}+i-2$, color $v_{i-1,j+1}$ with color 1 and all the remaining vertices with color 2(ri)+22(r-i)+2 to obtain a mutual-visibility coloring of $GT(r)$ with 2(ri)+22(r-i)+2 colors.
@@
Sub case 2: $2^{i-1}+i-1\leq r\leq 2^{i-1}+2^{i-2}+i-3$.
The coloring procedure is the same as Case 1, where the only difference is that more than half of the vertices of depth i1i-1 (all vertices when $r=2^{i-1}+i-1$) will be uncolored. We can use one color each for these uncolored vertices in $V_{r}^{(1)}$ and $V_{r}^{(2)}$ so that the resulting coloring is a mutual-visibility coloring of $GT(r)$ by Lemma 3.5.
To prove the converse, first we prove that 2(ri)+12(r-i)+1 colors are not enough for a mutual-visibility coloring of $GT(r)$, where $i$ is the unique integer such that $2^{i-1}+i-2\leq r\leq 2^{i}+i-2$. Assume the contrary that there exists a mutual-visibility coloring with 2(ri)+12(r-i)+1 colors.
@@
Now, consider the cycles $C_{t2^{r-2},t2^{r-2}+1}$, for $t=1,3$. These two cycles have 4r124r-12 vertices each which are not part of both the previous cycles considered. Similar to the case above we can find a set of colors $\chi_{2}$ which are used three times either in $V(C_{2^{r-2},2^{r-2}+1})\setminus\{V(C_{1,2^{r}})\cup V(C_{2^{r-1},2^{r-1}+1})\}$ and $V(C_{2^{r-1}+2^{r-2},2^{r-1}+2^{r-2}+1})\setminus\{V(C_{1,2^{r}})\cup V(C_{2^{r-1},2^{r-1}+1})\}$. Also, $\chi_{2}\cap\chi_{1}=\chi_{2}\cap\chi_{0}=\phi$ and $|\chi_{2}|\geq 2(4i-12)$. Now, we consider the four cycles $C_{t2^{r-3},t2^{r-3}+1}$, for $t=1,3,5,7$, and proceed as above to get $\chi_{3}$ of cardinality at least 4(4i16)4(4i-16) which do not intersect pairwise with any of $\chi_{0},\chi_{1}$ and $\chi_{2}$. Continue like this to get a collection of distinct colors $\chi_{1},\chi_{2},\ldots,\chi_{i-1}$, where each $\chi_{j}$ is of cardinality at least $2^{i-1}(4i-4j-4)$, $j\in[i-1]$ and $\chi_{0}$ of cardinality at least 4i24i-2. Note that the disjoint sets $\chi_{j}$, $j\in[i-1]$ have cardinality at least $2^{i-1}(4i-4j-4)$ which is an arithmetic-geometric progression with common difference -4 and common ratio 2. Therefore, $|\cup_{j=1}^{i-1}\chi_{j}|\geq 2^{i+1}-4i$ and together with the 4i24i-2 more distinct colors in $\chi_{0}$, at least $2^{i+1}-2$ colors are used three times to color $GT(r)$. But, we have only 2(ri)+12(r-i)+1 colors and $r\leq 2^{i}+i-2$ so that, $2(r-i)+1\leq 2^{i+1}-3$, a contradiction.
@@
The general position variation of mutual-visibility chromatic number was introduced by [4] in 2024. The minimum number of colors required to color the vertices of $G$ such that each color class is in general position is called general position chromatic number, denoted by $\chi_{\text{gp}}(G)$. The color classes obtained in the proof of Theorem 3.6 are also in general position except for $r=2^{i-1}+2^{i-2}+i-2$. Hence, we have the following corollary.
@@
Let $S$ be a mutual-visibility set of $GT(r,t)$. Let $C_{n,k}$ denote the circle passing through the quasi-leaves $v_{n}$ and $v_{k}$. Then $|{S\cap V(C_{i,j})}|\leq 3$.
@@
We proceed with the same coloring pattern used in Theorem 4.6 so that we reach a stage where the vertices of depth i1i-1 in $V_{r}^{(2)}$, $v^{\prime}_{i,1},v^{\prime}_{i,2},\ldots,v^{\prime}_{i,j}$ will be already colored, where $j=r-i-(\frac{t^{i-1}-1}{t-1})\geq\frac{t^{i-1}}{2}-1$, (since, $r\geq\frac{t^{i-1}-1}{t-1}+\frac{t^{i-1}}{2}+i-1$). Similarly, in $V_{r}^{(2)}$, $v^{\prime}_{i,a},v^{\prime}_{i,a-1},\ldots,v^{\prime}_{i,a-j-1}$ will be already colored, where $a=t^{i-1}$ and $j\geq\frac{t^{i-1}}{2}-1$. When $r\geq\frac{t^{i-1}-1}{t-1}+\frac{t^{i-1}}{2}+i-1$, we have $j\geq\frac{t^{i-1}}{2}$ so that there are fewer than $\frac{t^{i-1}}{2}$ uncolored vertices of depth i1i-1. In this case, color all the remaining uncolored vertices with color 2(ri)+22(r-i)+2. When $r=\frac{t^{i-1}-1}{t-1}+\frac{t^{i-1}}{2}+i-1$, color $v_{i-1,j+1}$ with color 1 and all the remaining vertices with color 2(ri)+22(r-i)+2 to obtain a mutual-visibility coloring of $GT(r,t)$ with 2(ri)+22(r-i)+2 colors.
@@
Sub case 2: $\frac{t^{i-1}-1}{t-1}+i\leq r\leq\frac{t^{i-1}-1}{t-1}+\frac{t^{i-1}}{2}+i-2$.
The coloring procedure is the same as Case 1, where the only difference is that more than half of the vertices of depth i1i-1( all vertices when $r=\frac{t^{i-1}-1}{t-1}+i$) will be uncolored. We can use one color each for these uncolored vertices in $V_{r}^{(1)}$ and $V_{r}^{(2)}$ so as to obtain a mutual-visibility coloring of $GT(r,t)$.
@@
Consider the cycle $C_{1,t^{r}}$. By Lemma 3.8, a color can be used at most 3 times to color the vertices of this cycle. Since we have only 2(ri)+12(r-i)+1 colors, and there are $4r$ vertices in $C_{1,t^{r}}$, at least 4r2[2(ri)+1]=4i24r-2[2(r-i)+1]=4i-2 colors must be used at least three times. Let $\chi_{0}$ be the collection of colors which are used three times to color the vertices of $C_{1,t^{r}}$ so that $|\chi_{0}|\geq 4i-2$. Now, consider (t1)(t-1) cycles $C_{mt^{r-1},mt^{r-1}+1}$ where $m\in[t-1]$, note that all the cycles have length $4r$. Let $S_{m}=V(C_{mt^{r-1},mt^{r-1}+1})\setminus\{V(C_{1,t^{r}})\cup V(C_{(m-1)t^{r-1},(m-1)t^{r-1}+1})\}$, $m\in[t-1]$ . Note that there are 4r44r-4 vertices in $S_{m}$, if $m\in[t-2]$ and 4r64r-6, if m=t1m=t-1. Similar to the discussion above, at least 4i64i-6 colors must be used three times to color those vertices in $S_{m}$, $m\in[t-2]$ and at least 4i84i-8 colors must be used three times to color vertices in $S_{t-1}$. Let $\chi_{m}$ denotes the collection of colors which are used three times to color the vertices in $S_{m}$ , $m\in[t-1]$ so that $|\chi_{m}|\geq 4i-6$, where $m\in[t-2]$ and $|\chi_{m}|\geq 4i-8$, when m=t1m=t-1.
@@
Generally in $n^{th}$ step we consider $(t-1)t^{n-1}$ cycles $C_{mt^{r-n},mt^{r-n}+1}$ where $m\in[t^{n}]-t\mathbb{N}$. Note that in the $n^{th}$ step all the cycles have length 4(rn+1)4(r-n+1). Set $S^{0}_{m}=V(C_{1,2^{r}})$ and define $S_{m}^{n}=V(C_{mt^{r-n},mt^{r-n}+1})\setminus\{V(C_{1,t^{r}})\cup V(C_{(m-1)t^{r-n},(m-1)t^{r-n}+1})\cup S_{m}^{n-1}\}$, for every $n\in[i-1]$ . Note that there are 4r4n4r-4n vertices in $S^{n}_{m}$, if $m\not\equiv 1\pmod{t}$ and 4r4n24r-4n-2, if $m\equiv 1\pmod{t}$. Similar to the discussion above, since we have only 2(ri)+12(r-i)+1 colors, at least 4r4n2[2(ri)+1]=4i4n24r-4n-2[2(r-i)+1]=4i-4n-2 colors must be used at least three times to color those vertices in $S_{m}^{n}$, where $m\not\equiv 1\pmod{t}$, and at least 4i4n44i-4n-4 colors must be used three times to color those vertices in $S_{m}^{n}$, where $m\equiv 1\pmod{t}$. Let $\chi^{n}_{m}$ denotes the collection of colors which are used three times to color the vertices in $S_{m}^{n}$, $m\in[t-1]-t\mathbb{N}$ so that $|\chi^{n}_{m}|\geq 4i-4n-2$, if $m\not\equiv 1\pmod{t}$ and $|\chi^{n}_{m}|\geq 4i-4n-4$, if $m\equiv 1\pmod{t}$.